# Acceleration without photon pair creation

### v1

@@ Appendix A General frame transformations
To use the results that we obtained in the main text when considering an observer moving at a constant speed $v_{\rm B}$, we approximate Bob’s trajectory in the following by a succession of short time intervals of constant motion. As illustrated in Fig. 3, we assume that Bob’s speed changes whenever he crosses a light-like trajectory (red lines), which has been highlighted by one of Alice’s explosions. In the limit where $\Delta\chi_{\rm A}$ tends to zero, Bob’s trajectory should converge to his actual trajectory given by $(t_{\rm A},\mu_{\rm A}(t_{\rm A}))$ in Alice’s spacetime diagram. Suppose $\chi_{\rm A}^{(N)}$ and $\chi_{\rm B}^{(N)}$ are the natural coordinates of the light line created by $n$-th explosion. Using this notation, we know that
@@
with $\Delta\chi_{\rm B}^{(n)}=\chi_{\rm B}^{(n)}-\chi_{\rm B}^{(n-1)}$. In Figs. 3 and 4 respectively, $\Delta\chi_{\rm A}$ and $\Delta\chi_{\rm B}^{(n)}$ coincide with the lengths of the bottom line of the light blue triangles.
@@
In the following, we denote the times Alice triggers the $n^{\text{th}}$ explosion by $t_{{\rm A}}^{(1,n)}$ and $t_{{\rm B}}^{(1,n)}$ respectively, while $t_{{\rm A}}^{(2,n)}$ and $t_{{\rm B}}^{(2,n)}$ mark the arrival times of the generated local light pulses at Bob’s position, and also the times when Bob changes his speed. Since Bob moves at a constant speed in each segment, we can assume that Alice’s and Bob’s clocks progress at a constant speed and distances transform linearly between two subsequent light lines. Without restrictions, we therefore know that $\Delta\chi_{\rm B}^{(n)}/\Delta\chi_{\rm A}=\kappa\big{(}\chi_{\rm A}^{(n)}\big{)}$ and that $\Delta t_{{\rm B}}^{(1,n)}/\Delta t_{{\rm A}}^{(1,n)}=\Delta t_{{\rm A}}^{(2,n)}/\Delta t_{{\rm B}}^{(2,n)}=\gamma\big{(}\chi_{\rm A}^{(n)}\big{)}$ with $\Delta t_{i}^{(1,n)}=t_{i}^{(1,n)}-t_{i}^{(1,n-1)}$, in analogy to the definition of the parameters $\kappa$ and $\gamma$ in the main text. Moreover, Figs. 3 and 4 suggest that
@@
Now we can relate $\chi_{\rm A}$ to $\chi_{\rm B}$ by taking the limit of small $\Delta\chi_{\rm A}$ to obtain Eq. (4) in the main text. For any given trajectory $(ct_{\rm A},\mu(t_{\rm A}))$ of Bob in Alice’s spacetime diagram, we can evaluate the relation between Alice’s and Bob’s natural coordinates. However, this only applies to Alice and Bob who have met at an initial time with $t_{\rm A}=t_{\rm B}=0$ and who remain standing at the origins of their respective $x$ axis.

### v2

@@ Appendix A General frame transformations
To use the results that we obtained in the main text when considering an observer moving at a constant speed $v_{\rm B}$, we approximate Bob’s trajectory in the following by a succession of short time intervals of constant motion. As illustrated in Fig. 3, we assume that Bob’s speed changes whenever he crosses a light-like trajectory (red lines), which has been generated by one of Alice’s explosions. In the limit where $\Delta\chi_{\rm A}$ tends to zero, Bob’s trajectory should converge to his actual trajectory given by $(t_{\rm A},\mu_{\rm A}(t_{\rm A}))$ in Alice’s spacetime diagram. Suppose $\chi_{\rm A}^{(N)}$ and $\chi_{\rm B}^{(N)}$ are the natural coordinates of the light line created by the $n$-th explosion. Using this notation, we know that
@@
with $\Delta\chi_{\rm B}^{(n)}=\chi_{\rm B}^{(n)}-\chi_{\rm B}^{(n-1)}$. In Figs. 3 and 4 respectively, $\Delta\chi_{\rm A}$ and $\Delta\chi_{\rm B}^{(n)}$ coincide with the lengths of the bottom edges of the light blue triangles.
@@
In the following, we denote the times Alice triggers the $n^{\text{th}}$ explosion by $t_{{\rm A}}^{(1,n)}$ and $t_{{\rm B}}^{(1,n)}$ in Alice’s and Bob’s frames respectively, while $t_{{\rm A}}^{(2,n)}$ and $t_{{\rm B}}^{(2,n)}$ mark the arrival times of the generated local light pulses at Bob’s position, and also the times when Bob changes his speed. Since Bob moves at a constant speed in each segment, we can assume that Alice’s and Bob’s clocks progress at a constant rate and distances transform linearly between two subsequent light lines. Without restrictions, we therefore know that $\Delta\chi_{\rm B}^{(n)}/\Delta\chi_{\rm A}=\kappa\big{(}\chi_{\rm A}^{(n)}\big{)}$ and that $\Delta t_{{\rm B}}^{(1,n)}/\Delta t_{{\rm A}}^{(1,n)}=\Delta t_{{\rm A}}^{(2,n)}/\Delta t_{{\rm B}}^{(2,n)}=\gamma\big{(}\chi_{\rm A}^{(n)}\big{)}$ with $\Delta t_{i}^{(1,n)}=t_{i}^{(1,n)}-t_{i}^{(1,n-1)}$, in analogy to the definition of the parameters $\kappa$ and $\gamma$ in the main text. Moreover, a closer look at Figs. 3 and 4 suggests that
@@
Now we can relate $\chi_{\rm A}$ to $\chi_{\rm B}$ by taking the limit of small $\Delta\chi_{\rm A}$ to obtain Eq. (4) in the main text. For any given trajectory $(ct_{\rm A},\mu(t_{\rm A}))$ of Bob in Alice’s spacetime diagram, we can now evaluate the relationship between Alice’s and Bob’s natural coordinates. However, this only applies to Alice and Bob who have met at an initial time with $t_{\rm A}=t_{\rm B}=0$ and who remain standing at the origins of their respective $x$ axis.